# Security Approaches for Data Provenance in the Internet of Things: A Systematic Literature Review

### v1

@@ Appendix A Additional Figures and Comparison Tables
Reference Provenance Category Implementation Simulation
 Lim et al. (2010) Cryptography-based Java
 Wang et al. (2016a) Cryptography-based TelosB motes TinyOS 2.1.2 TOSSIM
 Chia et al. (2017) Cryptography-based Raspberry Pi 3 and DLink DSP-W215 Wi-Fi Smart Plug Java
 Suhail et al. (2018) Cryptography-based Contiki simulator, Cooja  (Osterlind et al., 2006)
 Xu et al. (2019) Cryptography-based Testbed with Zigbee sensor nodes TinyOS 2.1.2/TOSSIM, PowerTOSSIMz
 Suhail et al. (2020) Cryptography-based Contiki simulator, Cooja
 Liu and Wu (2020) Cryptography-based Python
 Tang and Keoh (2020) Cryptography-based
 Porkodi and Kesavaraja (2021) Cryptography-based N/A
 Xu et al. (2022) Cryptography-based Zigbee sensor nodes with TI CC2530 microcontroller TOSSIM/TinyOS 2.1.2 and PowerTOSSIMz (Perla et al., 2008)
 Xu and Wang (2022) Cryptography-based ZigBee sensor/TI CC2530 microcontroller TinyOS 2.1.2/TOSSIM, PowerTOSSIMz
@@
 Siddiqui et al. (2019) Bloom Filter Java
 Harshan et al. (2020) Bloom Filters Raspberry Pi 3+ and a Digi XBee S2C
 Ali et al. (2014) Link Fingerprints MICAz motes, running TinyOS
 Alam and Fahmy (2014) Rabin Fingerprints TelosB motes TOSSIM
 Kamal and Tariq (2018) Link Fingerprints MICAz motes
 Baracaldo et al. (2017) Blockchain-based
 Zeng et al. (2018) Blockchain-based MICAz motes TOSSIM/TinyOS 2.1.2
 Javaid et al. (2018) PUFs and Blockchain Solidity IDE, Remix
 Sigwart et al. (2020) Blockchain-based Ethereum networks Rinkeby and Ropsten
 Liu et al. (2020) Blockchain-based Python interface/Pinocchio (Parno et al., 2016) and libsnark interface (Kosba et al., 2018) with Etheruem test network
 Siddiqui et al. (2020) Blockchain-based Hyperledger Fabric (Cachin, 2016)/ Byzantine fault-tolerant state machine replication library - Java (BFT-SMaRt) (Sousa et al., 2018)
 Yin and Fu (2022) Blockchain-based Ethereum network test/ Ubuntu
 Sun et al. (2022) Blockchain-based Public IoT database MMASH (Rossi et al., 2020)
 Chenli et al. (2022) Blockchain-based Go language and Hyperledger Fabric v2.0
@@
 Aman et al. (2019) PUFs and Link Fingerprints MICAz motes/ MATLAB MICA 2 mote platform/ AVRORA
@@
 Chang et al. (2022) Event-flow graphs Gentle Wake-Up, and Smart Security apps using SmartThings IDE (SmartThings, 2023; Co., 2023)
 Jaigirdar et al. (2023) Provenance graphs Flask module in python/ Neo4j
 Al-Rakhami and Al-Mashari (2022) Directed Acyclic Graph Eclipse Paho  [^2] https://www.eclipse.org/paho/ and Python
@@
 Lomotey et al. (2018) Lexical chaining techniques Bluetooth devices/Amazon EC2 platform
 Gao et al. (2013) Path difference TinyOS 2.1.1 with CTP (Gnawali et al., 2009)/ TOSSIM
@@
 Wang et al. (2018) Centralized Database Samsung SmartThings IDE (Samsung, 2017)
 Elkhodr and Alsinglawi (2020) Management platform
@@
 Lim et al. (2010) N/A N/A YES N/A YES N/A N/A
 Wang et al. (2016a) YES YES YES N/A YES YES N/A
 Chia et al. (2017) YES N/A N/A N/A N/A YES YES
 Suhail et al. (2018) YES N/A N/A N/A YES YES YES
 Xu et al. (2019) YES N/A N/A N/A N/A YES N/A
 Suhail et al. (2020) N/A N/A YES N/A YES YES N/A
 Liu and Wu (2020) YES N/A YES N/A YES N/A YES
 Tang and Keoh (2020) YES YES N/A N/A N/A N/A YES
 Porkodi and Kesavaraja (2021) YES YES YES N/A N/A N/A YES
 Xu et al. (2022) YES YES YES N/A N/A YES YES
 Xu and Wang (2022) YES N/A YES N/A N/A YES YES
 Sultana et al. (2015) YES YES YES N/A YES N/A N/A
 Siddiqui et al. (2019) YES YES YES N/A N/A YES N/A
 Harshan et al. (2020) YES N/A YES N/A N/A YES YES
 Ali et al. (2014) YES YES N/A N/A YES YES YES
 Alam and Fahmy (2014) N/A N/A N/A N/A N/A N/A N/A
 Kamal and Tariq (2018) YES N/A YES N/A YES N/A N/A
 Baracaldo et al. (2017) YES YES N/A YES N/A YES N/A
 Zeng et al. (2018) YES YES N/A N/A YES YES YES
 Javaid et al. (2018) YES N/A N/A N/A YES YES N/A
 Sigwart et al. (2020) YES N/A YES N/A N/A YES N/A
 Liu et al. (2020) YES N/A N/A N/A YES YES YES
 Siddiqui et al. (2020) YES YES N/A N/A YES YES N/A
 Yin and Fu (2022) YES YES N/A YES N/A YES N/A
 Sun et al. (2022) YES YES N/A N/A N/A N/A YES
 Chenli et al. (2022) YES YES N/A N/A N/A N/A N/A
 Sultana et al. (2011) YES YES YES N/A YES N/A N/A
 Aman et al. (2019) YES YES N/A YES YES YES N/A
 Aman et al. (2021) YES N/A N/A N/A YES N/A N/A
 Hamadeh and Tyagi (2021) YES N/A YES YES YES YES N/A
 Chang et al. (2022) N/A N/A N/A N/A YES N/A N/A
 Jaigirdar et al. (2023) YES N/A YES N/A N/A YES YES
 Al-Rakhami and Al-Mashari (2022) YES YES YES N/A N/A N/A YES
 Lomotey et al. (2019) N/A YES N/A YES N/A YES N/A
 Lomotey et al. (2018) YES YES YES N/A N/A YES N/A
 Gao et al. (2013) N/A N/A N/A N/A YES YES N/A
 Sadineni et al. (2023) N/A N/A YES N/A N/A N/A YES
 Alkhalil et al. (2019) N/A N/A YES N/A N/A YES N/A
 Wang et al. (2018) YES N/A YES YES N/A YES N/A
 Elkhodr and Alsinglawi (2020) N/A N/A N/A N/A YES N/A N/A

### v2

@@ Appendix A Additional Figures and Comparison Tables
Reference Provenance Implementation Simulation
Category Hardware Cloud-based Programming Language Simulator
Lim et al. (2010) Cryptography Java Program
Wang et al. (2016a) Cryptography TelosB motes TinyOS/TOSSIM
Chia et al. (2017) Cryptography Raspberry Pi 3/ DLink Smart Plug Java Program
Suhail et al. (2018) Cryptography Contiki simulator, Cooja  (Osterlind et al., 2006)
Xu et al. (2019) Cryptography Zigbee sensor nodes TinyOS/ PowerTOSSIMz
Suhail et al. (2020) Cryptography Contiki simulator, Cooja
Liu and Wu (2020) Cryptography Python Program
Tang and Keoh (2020) Cryptography
Xu et al. (2022) Cryptography Zigbee/ TI CC2530 microcontroller TinyOS PowerTOSSIMz (Perla et al., 2008)
Xu and Wang (2022) Cryptography ZigBee/ TI CC2530 TinyOS/ PowerTOSSIMz
@@
Siddiqui et al. (2019) Bloom Filter Java Program
Harshan et al. (2020) Bloom Filter Raspberry Pi 3+ and a Digi XBee S2C
Ali et al. (2014) Fingerprints MICAz motes, running TinyOS
Alam and Fahmy (2014) Fingerprints TelosB motes TOSSIM
Kamal and Tariq (2018) Fingerprints MICAz motes
Baracaldo et al. (2017) Blockchain
Zeng et al. (2018) Blockchain MICAz motes TOSSIM/TinyOS 2.1.2
Javaid et al. (2018) Blockchain Solidity IDE, Remix
Sigwart et al. (2020) Blockchain Ethereum (Rinkeby and Ropsten)
Liu et al. (2020) Blockchain Python interface/Pinocchio (Parno et al., 2016) and libsnark interface (Kosba et al., 2018) with Etheruem test network
Siddiqui et al. (2020) Blockchain Hyperledger Fabric (Cachin, 2016)/ Java (BFT-SMaRt) (Sousa et al., 2018)
Porkodi and Kesavaraja (2021) Blockchain N/A
Yin and Fu (2022) Blockchain Ethereum network test/ Ubuntu
Sun et al. (2022) Blockchain Public IoT database MMASH (Rossi et al., 2020)
Chenli et al. (2022) Blockchain Go language and Hyperledger Fabric v2.0
@@
Aman et al. (2019) PUFs MICAz motes/ MATLAB MICA 2 mote platform/ AVRORA
@@
Chang et al. (2022) Graph-based Gentle Wake-Up/ SmartThings IDE (SmartThings, 2023; Co., 2023)
Jaigirdar et al. (2023) Graph-based Flask module in python/ Neo4j
Al-Rakhami and Al-Mashari (2022) Graph-based Eclipse Paho  [^2] https://www.eclipse.org/paho/ and Python
@@
Lomotey et al. (2018) Lexical chain Bluetooth devices Amazon EC2 platform
Gao et al. (2013) Path difference TinyOS with CTP (Gnawali et al., 2009)/ TOSSIM
@@
Wang et al. (2018) Storing Samsung SmartThings IDE (Samsung, 2017)
Elkhodr and Alsinglawi (2020) Storing
@@
[…]
 Alam and Fahmy (2014)
[…]